ABC company is under contract to make 10,000 blenders. It must decide whether to make or buy the blender motor. Unfortunately, ABC does not presently have motor technology. Motor development is a two-stage process. The first stage has a 80% chance of success and will cost $30,000. The second stage has a 60% chance of success and will cost $20,000. If development is successful, then the variable cost of producing a motor will be $2.50/motor. If either development stage is unsuccessful, ABC must purchase motors outside at a cost of $10/motor. a) Should ABC purchase the motor outside or begin development? b) What is the most ABC would pay to know, before making the first decision, whether the first stage would succeed? the second stage?
Answer #1

solution ;

given that

ABC company is under contract to make 10,000 blenders.

also given that

It must decide whether to make or buy the blender motor. Unfortunately, ABC does not presently have motor technology

we know that

Motor development is a two-stage process.

a) Should ABC purchase the motor outside or begin development?

Make or Buy Decision relies upon which alternative is progressively conservative, which means costs less.

Cost of Buying = Total Blenders * Cost per Blender

= 10,000* 10 = $100,000

Cost of making = Sunk Cost + Opportunity Cost at each stage + Total Variable Cost

Sunk Cost = $30,000 + 20,000 = $50,000

Opportunity cost = Expected Cost of Failure at every one of improvement arrange

= (1-.8)*30,000 + (1-.6)*20,000 = 6000 + 8000 = $14,000

Complete Variable Cost = 10,000 * 2.5 = $25,000

Cost of making in-house = 50,000 + 14,000 + 25, 000 = $89,000

Since making cost is not as much as obtaining from outside, ABS ought to begin advancement.

b) What is the most ABC would pay to know, before making the first decision, whether the first stage would succeed? the second stage?

Let x be sum AMB willing to pay for knowing apriori whether Stage 1 would suceed.

Cost to firm after disappointment = Investment in Stage 1 + Cost of purchasing from outside

= 30,000 + 100,000 = $130,000

On the off chance that ABC chooses to purchase, cost is just 100,000

In this manner, 100,000 + x = 130,000 or x = $30,000

Applying a comparable rationale for Stage 2,

100,000 = 30,000 + 20,000 + y + 25,000

or on the other hand y = $25,000

The $25,000 originates from reality that in the event that Stage 2 likewise succeeds, cost to create blender is whole of all out interest being developed + VC of motorsa) Make or Buy Decision relies upon which alternative is progressively conservative, which means costs less.

Cost of Buying = Total Blenders * Cost per Blender

= 10,000* 10 = $100,000

Cost of making = Sunk Cost + Opportunity Cost at each stage + Total Variable Cost

Sunk Cost = $30,000 + 20,000 = $50,000

Opportunity cost = Expected Cost of Failure at every one of advancement arrange

= (1-.8)*30,000 + (1-.6)*20,000 = 6000 + 8000 = $14,000

All out Variable Cost = 10,000 * 2.5 = $25,000

Cost of making in-house = 50,000 + 14,000 + 25, 000 = $89,000

Since making cost is not as much as buying from outside, ABS ought to begin advancement.

b) Let x be sum AMB willing to pay for knowing apriori whether Stage 1 would suceed.

Cost to firm after disappointment = Investment in Stage 1 + Cost of purchasing from outside

= 30,000 + 100,000 = $130,000

In the event that ABC chooses to purchase, cost is just 100,000

In this manner, 100,000 + x = 130,000 or x = $30,000

Applying a comparable rationale for Stage 2,

100,000 = 30,000 + 20,000 + y + 25,000

or on the other hand y = $25,000

The $25,000 originates from reality that on the off chance that Stage 2 likewise succeeds,

therfore

cost to create blender is entirety of all out interest being developed + VC of engines